### 06 Mar 2026 v1

# LLM2SMT: Building an SMT Solver with Zero Human-Written Code

Abstract
Whether LLMs can reason or write software is widely debated, but whether they can write software that itself reasons is largely unexplored. We present a case study in which an LLM coding agent builds a complete DPLL(T)-style SMT solver for QF_UF with zero human-written code. The solver implements the Nieuwenhuis-Oliveras congruence closure algorithm, includes preprocessing, and emits Lean proofs for unsatisfiable instances. We describe the development process and key challenges, and show that the resulting solver is competitive on SMT-LIB benchmarks.
keywords:
LLM, SMT, automated reasoning, proofs
category:
[…]
1 Introduction
LLM-powered coding agents are now so widely used that some believe software engineers in their traditional role will disappear [27]. LLMs—directly or via coding agents—have also been used for auto-formalization and for developing novel proofs [18, 3, 31, 16]. The question we ask is: Can LLMs develop automated reasoning tools?
The appeal is clear: rapid feature integration, experimentation with new techniques, and importing methods from recent publications. Skepticism is warranted, however—small mistakes can propagate into subtle, hard-to-detect errors. Correctness is paramount: do LLMs understand the underlying logic well enough to write reliable SMT solvers?
We develop a basic DPLL(T)-style [25] SMT solver with zero code written by a human, focusing on the theory of uninterpreted functions without quantifiers (QF_UF)—an important theory underlying most functionality in DPLL(T).
The solver uses CaDiCaL [8] via the IPASIR-UP interface [15], emits proofs in Lean [21] (under some limitations), implements the Nieuwenhuis–Oliveras congruence closure algorithm [24], and integrates several preprocessing techniques. The code is available as a GitHub repository [^1] https://github.com/MikolasJanota/llm2smt.
2 Background
We assume that the reader is familiar with classical first-order logic with equality [29].
Satisfiability Modulo Theories (SMT) solving is a well-established research area encompassing a wide range of logical theories, applications, and algorithmic techniques. Most SMT solvers reduce the input formula to propositional satisfiability (SAT), either eagerly [26], where theory reasoning is encoded directly into a propositional (Boolean) problem, or lazily [10], where Boolean reasoning is interleaved with theory-specific checks. These checks result in theory lemmas, which strengthen the propositional problem.
Most modern solvers are based on the DPLL (T) framework [25, 17, 28]. In this architecture, theory atoms are abstracted as Boolean variables and handled using conventional SAT-solving techniques [19]. Candidate assignments produced by the SAT solver are subsequently validated by dedicated theory solvers, which are required only to check conjunctions of theory literals. If a conjunction is inconsistent in the underlying theory, the theory solver provides conflict information to refine the Boolean search. The process iterates until the SAT solver determines the formula to be unsatisfiable or a theory-consistent assignment is obtained. During the search, the solver might employ theory propagation, forcing certain literals to specific values under the current partial assignment.
This paper focuses on the theory of equality with uninterpreted functions (QF-EUF), which is the basis of equality reasoning; in particular, it is instrumental in the Nelson-Oppen closure [22, 11], which enables cooperation between separate theory solvers under appropriate assumptions. In the context of SMT, quantifiers are handled by quantifier instantiation, cf. [14, 20]. Therefore, solving problems without quantifiers (ground problems) is necessary for solving quantified problems.
Lean [21] is a popular modern interactive theorem prover (ITP) and programming language for formalizing mathematics and verifying software. It is based on dependent type theory, providing a unified framework for programs, proofs, and mathematical objects. Its main library, mathlib, contains thousands of formalized theorems spanning broad areas of mathematics.
3 Case Study
The entire study was conducted within Claude Code using the Sonnet 4.6 model [2]. No code was written by a human, and the code was only minimally inspected. A human performed more extensive evaluation on a separate server.
The project was started with a minimal description giving the LLM a reference for the congruence closure algorithm [24].
Start a project that implements an SMT solver. We need a parser, a SAT solver, for which we will use the ipasir-up interface hash-consing of formulas. Then theory specific things, like arbitrary precision library. Setup build by using cmake. The solver will be written in c++20. Use antlr for parsing (the grammar is here https://github.com/julianthome/smtlibv2-grammar). Set up the directory structure also testing. cmake should download the necessary dependencies as needed.
We will start with QF-EUF theory solver implementing the congruence closure as described in Fast congruence closure and extensions by Robert Nieuwenhuis and Albert Oliveras (the pdf is in the folder for convenience).
This quickly proved too naive. Despite correctly describing SMT solving at a high level, the LLM did not implement a correct solver from the start. Most strikingly, the solver did not handle Boolean connectives at all. This was quickly remedied by pointing the LLM to the core theory specification [^2] https://smt-lib.org/theories-Core.shtml. Only after some time did we notice that xor was not implemented as n-ary, which required an explicit prompt to fix.
Despite being prompted to use CaDiCaL, the agent instead wrote its own rudimentary SAT solver. We subsequently gave an explicit prompt to integrate CaDiCaL via the IPASIR-UP interface [15], which the agent accomplished without difficulty.
Notably, the agent ran the solver without a timeout, leaving long-running processes unnoticed in the background. We therefore added to the permanent instructions of the LLM to always run the solver under the timeout command. [^3] Claude Code permits inserting such instructions into CLAUDE.md.
3.1 Bugs and Debugging
Initial development required numerous bug fixes in the congruence closure algorithm and parsing. The agent fixed bugs autonomously, once presented with an input file where the solver gave a wrong answer. Several issues stemmed from the fact that in SMT2, unlike in first-order logic, there is no distinction between formulas and terms (for example, $f(\forall x.\,x>0)=3$ is valid in SMT2 as long as $f$ is from Bool to Int). The following is a commit message written by the LLM for one of the bug fixes.
link_bool_term_to_euf: add mutual-exclusivity clause -eq_true, -eq_false so the SAT solver cannot simultaneously merge __bool_true and __bool_false, which was causing false UNSAT for formulas with Bool-sorted UF arguments.
The model handles bug fixing well, but—like human developers—benefits from structured debugging support: fuzzing [30] and delta-debugging [32]. We had the LLM generate a random formula generator and a differential testing script comparing our solver against a reference solver, rather than using existing SMT fuzzers [9]. These techniques also reduce token usage: without them, the LLM would pinpoint bugs “manually”—correct but expensive.
3.2 Preprocessing
SMT solvers typically have a preprocessing module for simplifying common benchmark constructs. We instructed the agent to handle unit propagation as well as short-circuit simplifications of conjunction and disjunction. Interestingly, a cursory manual inspection identified that trivial simplifications were missing, such as that $t=t$ simplifies to true and $\texttt{if }C\texttt{ then }t\texttt{ else }t$ simplifies to $t$.
An interesting set of benchmarks contained in the SMT-LIB are the equational diamond problems of the following form.
$(x_{i}=z_{i}\land z_{i}=x_{i+1})\lor(x_{i}=v_{i}\land v_{i}=x_{i+1}),\ i\in 1..n$ (1)
$x_{1}\neq x_{n+1}$ (2)
The problems are trivially unsatisfiable but difficult for DPLL(T) because it cannot infer literals not already present in the formula. Here, it would be useful to have access to the literal $x_{i}=x_{i+1}$, which trivially follows from the disjunct (1) and would quickly give us a contradiction with (2). Not having access to this literal, the solver is forced to explore $2^{n}$ combinations, which produce one theory lemma each. One could explore more advanced techniques to handle such issues internally [7, 23] but it is also a good candidate for preprocessing. We prompted the agent with a simple prompt.
Consider SMT problem with disjunctions linking a sequence of $x_{i}$ as $x_{i}=y_{i}=x_{i+1}$ or $x_{i}=z_{i}=x_{i+1}$. With $x_{1}\neq x_{100}$. This causes issues for lazy SMT and problems of this type are in QF_UF/eq_diamond. Devise preprocessing technique to overcome the exponential behavior on this.
From this prompt alone, the agent devised a preprocessing technique that solves the diamond problems instantaneously. For each disjunction, it computes the EUF-closure of equalities in each branch and extracts those common to all branches. These common consequences are added as new unit equality formulas to the formula list. For example, for $(x=y\land y=z)\lor(x=w\land w=z)$ branch 1 gives the closure: $\{x,y,z\}$, branch 2 gives the closure: $\{x,w,z\}$ with shared nodes: $\{x,z\}$ and common pair $x=z$.
3.3 Certification
For a satisfiable input formula, the produced model is validated by encoding it back as an SMT problem and checking it with a reference solver. We asked the agent to integrate such a check into our fuzzing script. The prompt was bundled with the specification of the get-model command, for which we used a part of the SMT2 standard.
Implement the get-model command, see get-model-spec.pdf for specification create a script to test a model, i.e., given an SMT2 file, run our solver, if SAT, add the model into the file and call (another) solver to check it’s indeed SAT.
This turned out to be quite easy for the agent. It initially had a bug that it did not introduce explicit constants to represent the members of the (finite) universe, but it realized that quickly on its own. Another subtle error arose during model construction on formulas of the form f(distinct(c0,c2))—again due to Booleans being treated both as propositions and as terms.
For unsatisfiable input formulas, we ask the agent to produce a Lean proof. More advanced and more suitable approaches to proof generation for SMT have been investigated in the literature [4, 6, 1] but we were mainly interested in whether the agent could formally capture the reasoning behind its own code.
Since we only rely on QF-EUF, all the relevant primitives readily translate to Lean primitives, e.g., equality in SMT translates to equality in Lean. We do not build a low-level Lean proof; instead, we rely on two automation tools:
  • grind to verify theory lemmas – clauses following directly from congruence closure.
  • bv_decide to verify the propositional fragment. In order to use bv_decide, we asked the agent to wrap each proposition into decide (...)
Both tools are more powerful than strictly needed, but neither alone can solve standard SMT problems without extra hints from the solver.
Initially, all types, constants, and SMT assumptions were placed as Lean hypotheses within a single theorem per SMT problem, with theory lemmas as have statements. This proved infeasible: some SMT problems have thousands of assumptions, and Lean struggled to parse even the problem alone. We instead adopted a less ITP-centric design: SMT types, constants, and assumptions become Lean axioms; each theory lemma is a separate theorem (discharged by grind); and the final False proof collects all of these via have before closing with bv_decide.
In particular, for the equational diamond problems (see Section˜3.2), the agent struggled to grasp the distinction between theory lemmas and the final proof. It was trying to close propositionally nontrivial subgoals with grind, which did not work. Eventually, we had to write down our own example of a Lean proof correctly handling equational diamonds. Then it was able to incorporate such proofs into its proof export.
Listing 1: Example of a final version of Lean export with equational diamonds, passed Lean verification. Note that the original statement has only one assert, and the proof only one theory lemma (except for preprocessing), otherwise, we would have multiple axioms, or multiple theorems.
import Std.Tactic.BVDecide
noncomputable section
open Classical
axiom U : Type
axiom x0 : U
axiom y0 : U
axiom z0 : U
axiom x1 : U
axiom y1 : U
axiom z1 : U
axiom x2 : U
axiom y2 : U
axiom z2 : U
axiom hyp1 : (((decide (x0 = y0) ∧ decide (y0 = x1)) ∨ (decide (x0 = z0) ∧ decide (x1 = z0))) ∧ ((decide (x1 = y1) ∧ decide (y1 = x2)) ∨ (decide (x1 = z1) ∧ decide (x2 = z1))) ∧ $\neg$(decide (x0 = x2)))
theorem eq_bridge_0 : decide (x0 = y0) → decide (y0 = x1) → decide (x0 = x1) := by grind
theorem eq_bridge_1 : decide (x0 = z0) → decide (x1 = z0) → decide (x0 = x1) := by grind
theorem eq_bridge_2 : decide (x1 = y1) → decide (y1 = x2) → decide (x1 = x2) := by grind
theorem eq_bridge_3 : decide (x1 = z1) → decide (x2 = z1) → decide (x1 = x2) := by grind
theorem cl_0 : decide (x0 = x1) → decide (x1 = x2) → decide (x0 = x2) := by grind
theorem contradiction : False := by
have hyp1 := hyp1
have eq_bridge_0 := eq_bridge_0
have eq_bridge_1 := eq_bridge_1
have eq_bridge_2 := eq_bridge_2
have eq_bridge_3 := eq_bridge_3
have cl_0 := cl_0
bv_decide
The correctness of the proof mechanism was probably the most challenging task of this case study. Repeated fuzzing and bug fixing were required for the agent to correctly implement all steps of the proof. Another obstacle is that the agent does not grasp the complexity at the Lean proof level. We had trouble communicating to the agent that only easy steps should be presented to the grind tactic. Table˜2, in the upcoming section, presents the experimental results on the whole benchmark set.
4 Performance
We evaluated the developed solver on QF-UF non-incremental benchmarks available in SMT-LIB [^4] https://zenodo.org/records/1688772. We also compare to cvc5 [5] and z3 [12]. The experiments were performed on a machine with two AMD EPYC 7513 32-Core processors and with 514 GiB RAM with 100 problem instances run in parallel with 180 s timeout.
Table 1: Number of solved instances. A win is counted if the solver is not worse than the best solver by 1s.
solver solved wins
z3 7,500 7,491
cvc5 7,494 7,308
llm2smt 7,468 6,486
llm2smt-no-prepro 7,355 6,150
llm2smt-no-th-prop 7,478 6,593
Figure 1: Comparison with and without theory propagation.
Number of instances W. Preprocessing W/o Preprocessing
Certified 285 259
Grind Failed 2,244 1,949
Timeout 1,314 1,472
Maximum Heartbeats 240 247
Stack Overflow 86 85
Table 2: Overview of the Certification Experiments
We compare our solver against z3 [12] and cvc5 [5], as well as ablations with preprocessing and theory propagation disabled. Table˜1 compares the number of instances solved and gives some indication of speed—for each solver a win is counted if it is not worse than the best time by 1s. Interestingly, the version without theory propagation performs slightly better than the one with. This is likely because propagation introduces overhead without payoff on these benchmarks. Figure˜1 compares the runtime of the version with preprocessing and the version without, excluding problems solved in under 1 second. The plot confirms that turning off theory propagation actually makes the solver faster. However, there are problems where it does make a considerable difference. This suggests that smarter propagation scheduling could help—for instance, problems with few function symbols less likely to benefit from it.
Table˜2 summarizes the certification experiments. Lean has its own limits on resources. Heartbeats measured the overall time, the grind tactic is limited on how many times it splits. Additionally, there is a limit on recursion stack. Further, it is not particularly the appropriate tool for this task and simply times out on some problems. This has resulted in a large number of failed certification attempts. Nevertheless, no erroneous proofs were found.
5 Observations and Recommendations
We conclude with a list of observations and recommendations that hopefully will be useful to other researchers pursuing a similar goal.
  • The input language must be very carefully specified (SMT2 in our case).
  • Tools that may be non-terminating or exhibit exponential behavior must be run under explicit resource limits.
  • The agent can debug its own code effectively when given standard tools: fuzzing and delta-debugging.
  • Have the agent commit frequently and produce a regression test for each bug fix; it can then use git history to verify and bisect changes.
  • Despite high overall capability, the agent may fail unexpectedly on trivial tasks—e.g., the preprocessor initially did not simplify $s=s$. This phenomenon is known as jagged intelligence [^5] https://x.com/karpathy/status/1816531576228053133?s=20 [13] and represents a significant challenge for reasoning tools development.
  • Concrete examples and tests work best to improve the agent’s performance and reliability.
  • Proof-generation capabilities appeared to be rather hard to achieve for the agent.
6 Conclusion and Future Work
The central question we posed was: Can LLMs develop automated reasoning tools? The answer is a qualified yes. With appropriate scaffolding—systematic fuzzing, delta-debugging, explicit resource limits, and careful specification of input formats—a coding agent can produce a competitive SMT solver from scratch, solving nearly as many benchmarks as mature solvers such as z3 and cvc5—admittedly on a single theory. The agent is capable of implementing techniques described in scientific articles and debugging its own code when given concrete failing examples.
Two important caveats emerged. First, correctness cannot be assumed: the agent requires guidance, concrete examples, and systematic testing throughout; subtle errors (such as mishandling Booleans as both propositions and terms) can persist unnoticed without a fuzzer. Second, proof generation proved to be the hardest challenge: the agent must understand not just the solver logic but also the expectations of the proof checker (Lean in our case), which required significant human guidance.
In the future we aim to add more functionality to the solver but also prefer more careful analysis of the written code to suggest improvements to the agent.